Use the following functions that show quantity demanded and quantity supplied to answer the questions

Qd=600 - 30P

Qs=-300 + 120P

What are the equilibrium quantity


1
Expert's answer
2021-05-31T15:38:59-0400

At equilibrium:

"Q_d=Q_s \\

600-30P=-300+120P \\

900=150P \\

P= \frac{900}{150}=6 \\

Q_{eq}=600- 30 \times 6 = 600-180 = 420"

Answer: 420